OpenStudy (anonymous):

Determine whether the equation defines y as a function of x. x + y = 9

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

Can someone help me?

OpenStudy (anonymous):

I think y is not a function of x

OpenStudy (anonymous):

y is the function of x if you put a value in x you will get a unique value for y

OpenStudy (anonymous):

x2 + y2 = 1

OpenStudy (anonymous):

x^2+y^2=1 y is not the function of x

OpenStudy (anonymous):

x^2+y^2=1 y is not the function of x

OpenStudy (anonymous):

this is not ^2+y^2=1 its x + y = 9 which is a function

OpenStudy (anonymous):

(-2, -7), (3, -5), (6, -4), (9, -6), (10, -1)} this is a function